OpenStudy (anonymous):

Could someone help me learn how to do this?? I will give a medal and fan to whoever help me learn the best! What is the product of sqrt(3) (2b sqrt(49) - b)

OpenStudy (anonymous):

\[\sqrt{3}*(2b \sqrt{49} -b)\] is this what it looks like?

OpenStudy (anonymous):

Thats it!

OpenStudy (anonymous):

@robz8

OpenStudy (anonymous):

well we know the square root of 49 is 7. so you get \[\sqrt{3} * (2b*7-b)\] which would be \[\sqrt{3} * (14b - b)\] keep reducing\[\sqrt{3} *(13b)\] and you get \[13b \sqrt{3} \]

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

you're welcome!